OpenStudy (anonymous):

The Sugar Sweet Company is going to transport its sugar to market. It will cost $4375 to rent trucks, and it will cost an additional $175 for each ton of sugar transported. Let C represent the total cost (in dollars), and let S represent the amount of sugar (in tons) transported. Write an equation relating C to S I'm gonna have to plot this equation by the way...anyone able to help?

OpenStudy (anonymous):

Hi Tajar :) thanks for looking at my question

OpenStudy (anonymous):

So how much would it cost to ship 1 ton of sugar?

OpenStudy (anonymous):

175? or do you mean the total? 4550 then

OpenStudy (anonymous):

Hi...no problem

OpenStudy (anonymous):

wait so then what next?

OpenStudy (anonymous):

Correct...so C = 4375 + (175 x1)

OpenStudy (anonymous):

sorry..my computer keeps logging me in and out

OpenStudy (anonymous):

ohhhhh gotcha...technology can be so dumb

OpenStudy (anonymous):

To ship 2 tons of sugar, how much would that be?

OpenStudy (anonymous):

C= 4375 + (175 x 2)

OpenStudy (anonymous):

so if S=tons of sugar, what would the equation be?

OpenStudy (anonymous):

4375+175s

OpenStudy (anonymous):

Yes..good job!

OpenStudy (anonymous):

right, but it won't let me plot that...I'm doing a review and it won't let me plot that...

OpenStudy (anonymous):

C=4375 + 175S (S, C) Just plug in values (1, 4550) (2, 4725) (3, 5075)

OpenStudy (anonymous):

all three of those? can you tell me why? I want to learn this stuff, not just get the answer :)

OpenStudy (anonymous):

For the first pair (S, C), like we discussed, S=1 ton and you told me then C=$4550

OpenStudy (anonymous):

right

OpenStudy (anonymous):

For the second point, we could choose 10 tons which is (10, C)...what would C be?

OpenStudy (anonymous):

ohhhh you're amplifying it...I see

OpenStudy (anonymous):

So you can pick any points you want to plug in.

OpenStudy (anonymous):

can you check to see if I did this one right?

OpenStudy (anonymous):

(2x-3)(7x+1) 14x^2 + 2x - 21x - 4 14x^2 - 19x - 4

OpenStudy (anonymous):

I don't think i can simplify anymore right?

OpenStudy (anonymous):

that's right

OpenStudy (anonymous):

YAY

OpenStudy (anonymous):

14x^2 + 2x-21x-3 then you combined the x terms to get 14x^2-19x-3 Good job!

OpenStudy (anonymous):

sorry if this all seems random...I'm doing a review pretest before I take my actual test

OpenStudy (anonymous):

no worries-good luck!